OpenStudy (anonymous):

Given: csc θ = −(17/15), where 270° ≤ θ ≤ 360° and cot β = −(3/4) where 90° ≤ β ≤ 180°. Find the exact value of sin(θ + β). (3)

OpenStudy (dumbcow):

\[\sin(a+b) = \sin a \cos b + \sin b \cos a\] set up triangles to obtain sina,sinb,cos a, cos b  solve for missing sides of triangles to get needed trig ratios plug values into formula for sin(a+b)